Suppose there is a 69 percent chance that there will be traffic congestion on Bridge A, a 57 percent chance there will be traffic congestion on bridge B, and a 45 percent chance there will be traffic on both bridges. what is the probability that there will be traffic congestion on AT LEAST one bridge?
Answer by stanbon(75887)

Suppose there is a 69 percent chance that there will be traffic congestion on Bridge A, a 57 percent chance there will be traffic congestion on bridge B, and a 45 percent chance there will be traffic on both bridges. what is the probability that there will be traffic congestion on AT LEAST one bridge?
--------
P(A jam) = 0.69
P(B jam) = 0.57
P(A and B) = 0.45
---
P(jam on at least one) = 1 - P(jam on none)
P(jam on none) = P(A' and B')
----
P(A' and B') = P(A') + P(B') - P(A' or B')
P(A' or B') = 1-P(A and B)
-----
So P(A' and B') = 0.31 + 0.43 -[1-0.45]
= 0.74 - 0.55
= 0.19
------------
Therefore P(jam on at least one) = 1-0.19 = 0.81
=======================================================
cheers,
Stan H.
=========================
P(jam on at least one) = 1 - P(jam on none)